Sharp Company manufactures a product for which the following standards have been set: Standard Quantity or Hours Standard Price or Rate Standard Cost Direct materials 3 feet $ 5 per foot $ 15 Direct labor ? hours ? per hour ? During March, the company purchased direct materials at a cost of $45,210, all of which were used in the production of 2,500 units of product. In addition, 4,100 direct labor-hours were worked on the product during the month. The cost of this labor time was $30,750. The following variances have been computed for the month: Materials quantity variance $ 3,600 U Labor spending variance $ 2,750 U Labor efficiency variance $ 700 U Required: 1. For direct materials: a. Compute the actual cost per foot of materials for March. b. Compute the price variance and the spending variance. 2. For direct labor: a. Compute the standard direct labor rate per hour. b. Compute the standard hours allowed for the month’s production. c. Compute the standard hours allowed per unit of product.
Answer #1
Materials Quantity variance
(AQ used - Std Qty allowed)*SR
(x - 2500*3)*5 =3600U
5x - 37500 = 3600U
x =(3600+37500)/5
8220
a) actual cost = 45210/8220
5.5 per foot answer
(
b) Materials price variance
(Actual rate - standard rate)*AQ purchased
(5.5-5)*8220
4110 U answer
Spending variance = 3600U+4110U
7710 U answer
2a) labor spending varian 2,750 U
labor efficiency varian 700 U
labor price variance 2050 U
labor rate variance = ( Actual rate - standard rate)*Actual hours
(30750- x*4100) = 2,050
x = (30750-2050)/4100
7
Standard hours 7 per hour
b) Standard hours =
labor efficiency variance
(Actual hours - standard hours allowed)*Standard rate
(4100 -x ) *7 = 700
28700 - 7x = 700
x                       = (28700-700)/7
4000
Standard hours allowed 4,000 hours
c) Standard hours allowed per unit of product
4000/2500
1.6
Standard hours allowed per unit of product 1.6 hours per unit